OpenStudy (anonymous):

-2x +y = 0 5x + 3y = -11 Can someone help me and show me the steps to solve each system by substitution?

OpenStudy (anonymous):

-2x + y = 0 5x + 3y = -11 notice how in -2x + y = 0 when you move y to the right you get: -2x = -y divide both sides by -1 and you get y = 2x also notice how there is a "y" in 5x + 3y = -11 substitude the y in the second equation with the first and you get: 5x + 3(2x) = -11 5x + 6x = -11 11x = -11 x = -1 you can do the same with y.

OpenStudy (sweetburger):

make -2x+y=0 into y=+2x and then input this into the next equation for y and that becomes 5x+3(2x)=-11 so solve that and it becomes 5x+6x=-11 than 11x=-11 and that becomes x=-1 then plug x=-1 into the original equation so it becomes -2(-1)+y=0 and that becomes y=2

OpenStudy (sweetburger):

i solved for both y and x :P

OpenStudy (anonymous):

i just solved for x ^^;

OpenStudy (anonymous):

Now I just need to make sense of it. Thanks!